Hi I am a mother of a pre-algebra student and I am having a problem helping her with this word problem.
You are given $100 and need to buy cats, dogs and mice which also must equal 100. Cats cost $1.00, Dogs cost $15.00 and mice cost $.25. All the money must be gone and you have to use each animal at least once.
I would appreciate any help that you can give me.
Thank you
Stephanie Kilburn

Answer by askmemath(368)

Since you have to buy 1 of each animal,let's say you buy X Cats, Y Dogs and Z Mice.
Now the sum total of these animals is 100
i.e. X+Y+Z = 100
Now since each cat costs $1, X cats will cost 1x X = 1X
Similarly each dog will cost you 15Y
and Each mouse will cost 0.25Z
The sum total is $100
So
1X+15Y + 0.25Z = 100 (1)
X+Y+Z = 100 (2)
Can you solve from here? Subtracting (2) from (1) will give you an equation in terms of Y and Z. Solve for either variable and put it back into any equation to get your answer.